sovle. check your solution. -7=9+2g
solve. check your solution. 5h+2-h=22

Answer by AnlytcPhil(1806)

                          -7 = 9 + 2g

To get used to algebra, at first make sure every term has a sign either a + or
a - on its left side.  The -7 already has a sign on its left.  The 9 does not
have a sign left of it so we put a + left of it

                          -7 = +9 + 2g 

There is one term on the left, -7, and two terms on the right, +9 and +2g. 

There are no terms that will combine on the left or right.

1. We want to get rid of all terms on the left which DON'T contain the unknown
letter g.

2. 1. We want to get rid of all terms on the right which DO contain the unknown
letter g.

         The -7 does not contain a g and it is on the left side, so we need to
get rid of it by adding its opposite, +7 to both sides

                          -7 = +9 + 2g
                          +7   +7
                        ------------------

Add vertically:

                          -7 =  +9 + 2g
                          +7    +7
                        ------------------
                           0 = +16 + 2g

Now e have a new equation:

                           0 = +16 + 2g

The first thing we come to is the +16. It does not contain h and it is on the
right side, so we leave it alone.

The next thing we come to is the +2g on the right side. It contains a g so we
must get rid of the +2g on the right by adding its opposite -2g to both sides

                           0 = +16 + 2g
                      -2g          - 2g 
                     --------------------

Add vertically

                           0 = +16 + 2g
                     -2g           - 2g 
                     --------------------
                     -2g + 0 = +16 +  0

We can eliminate the zeros

                         -2g = +16

When we get it down to just one term on the left with the letter and just one
term on the right with just a number, that's when we DIVIDE both sides by
the coefficient of the unknown letter on the left:

                          -2g    +16
                         ————— = ————
                           -2     -2

Now you can cancel the -2's on the left and divide +16 by -2, getting -8 on the
right                             
                           1      -8
                          -2g    +16
                         ————— = ————
                           -2     -2                        
                            1      1
and get

                             g = -8 
____________________________________________________________________________
¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯

solve. check your solution.

                    5h + 2 - h = 22

The terms on the left are: +5h, +2, and -h.  There is only one term on the
right, namely +22.  Write the 5h as +5h, -h as -1h, and the 22 as +22

                  +5h + 2 - 1h = +22

The terms +5h and the term -1h combine as +4h since +5 and -1 give +4.

                       +4h + 2 = +22

The first term we come to is the +4h. It contains h and it is on the left, so
we leave it as is.

The next term is the +2.  It does not contain h and is on the left side so we
must get rid of it by adding its opposite, -2, to both sides:

                       +4h + 2 = +22
                            -2 =  -2
                     —————————————————
                       +4h + 0 = +20

Get rid of the 0
                           +4h = +20

When we get it down to just one term on the left with the letter and just one
term on the right with just a number, that's when we DIVIDE both sides by
the coefficient of the unknown letter on the left:

                          +4h    +20
                         ————— = ————
                           +4     +4

Now you can cancel the +4's on the left and divide +20 by +4, getting +5 on the
right                             
                           1      +5
                          +4g    +20
                         ————— = ————
                           +4     +4                        
                            1      1
and get

                             g = +5 

Edwin
AnlytcPhil@aol.com